# Estimating non-linear functionals of trawl processes

### v2

@@ Appendix A Computing the AVAR
$+\sum_{k=l^{\prime}}^{j-1}\chi_{k}\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,j,l}^{(1)}\chi_{j})$
@@
We analyze now separately the terms
@@
We often use the fact that that $\Delta_{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert$ is uniformly bounded (this can be seen as in the proof of Lemma 6). Furthermore, we are going to verify that
@@
where can be written as $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\iota=1}^{\mathrm{N}}\mathcal{S}_{l,l^{\prime}}^{n,\iota}$, $\mathrm{N}\in\mathbb{N}$ fixed, and each summand satisfies either
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}\upsilon_{n},$ (68)
uniformly on $l,l^{\prime}$ and with with $\upsilon_{n}=\mathrm{o}(n\Delta_{n})$; or
@@
for some constant $C\geq 0$. Note that these inequalities will imply the desired result. Indeed, by Jensen’s inequality we have that
@@
$\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert\rvert\overline{c}_{l^{\prime}}^{n}\rvert\frac{1}{(n\Delta_{n})^{2}}\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim\frac{1}{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}+1/n\Delta_{n}.$
In the latter case we further have that as $n\rightarrow\infty$
@@
    If $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1},$ by the continuity of $\dot{g}$
    $\frac{1}{n}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}\lesssim T/n\Delta_{n}\rightarrow 0.$
@@
    $\frac{1}{n}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}\lesssim\mathrm{O}(1/n\Delta_{n})+\frac{m_{n}}{n}\times\frac{1}{m_{n}\Delta_{n}}\int_{0}^{m_{n}\Delta_{n}}a(s)^{p+1}s\mathrm{d}s\rightarrow 0,$
@@
Therefore, in any situation, it holds that $\mathbb{E}(\rvert\mathcal{V}_{n}^{(\ell,\ell^{\prime})}-\mathbb{E}(\mathcal{V}_{n}^{(\ell,\ell^{\prime})})\rvert^{2})\rightarrow 0$, as $n\rightarrow\infty$, which is exactly the conclusion of this Lemma.
For convenience, when $\ell=\ell^{\prime}$, we will refer to $\mathcal{V}_{n}^{(\ell,\ell^{\prime})}$ as the conditional variance, otherwise we will use the adjective conditional covariance. We would also like to emphasize that when we present the decomposition (67) we are always going to use the notation $\mathcal{S}_{l,l^{\prime}}^{n}$ and its definition will vary from case to case.
@@
$\mathcal{S}_{l,l^{\prime}}^{n}:=$ $\mathfrak{K}_{3}\int_{t_{l}}^{t_{l+1}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l^{\prime}}^{(2)}$
$+\mathfrak{K}_{3}\int_{t_{l}}^{t_{l+1}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l}^{(2)}$
$+\int_{0}^{\Delta_{n}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}[\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)})]$
$=:$ $\sum_{\iota=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\iota},$
so that (67) holds. Since the sequence $\beta_{j,j-1,l^{\prime}}^{(2)}$ is $l^{\prime}$-dependent and $\mathbb{E}[(\beta_{j,j-1,l^{\prime}}^{(2)})^{2}]\leq\mathrm{Leb}(A),$ we have that
$\mathbb{E}\left[\left(\sum_{j=l\lor l^{\prime}}^{n}\beta_{j,j-1,l^{\prime}}^{(2)}\right)^{2}\right]\leq Cnl^{\prime},$
which adds up to $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}(l^{\prime}\lor l\Delta_{n}+C).$ Similar arguments can be applied ($\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}$ is $l\lor l^{\prime}$-dependent) to show that the same estimate holds for $\mathcal{S}_{l,l^{\prime}}^{n,2}$ and $\mathcal{S}_{l,l^{\prime}}^{n,3}$, which is enough.
@@
$\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(2)}\xi_{j,l^{\prime},n}^{(2)}\mid\mathscr{G}_{j-1}^{n})=\sum_{k=l}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{n-1}\chi_{k}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k\lor k^{\prime}+1}^{n-1}\mathcal{P}_{A}^{n}(i,j)).$
@@
$\mathcal{S}_{l,l^{\prime}}^{n}:=$ $\sum_{k=l}^{n-1}[\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]b_{k,l,l^{\prime}}+\sum_{k=l}^{n-1}\chi_{k}\nu_{k,l^{\prime},l}$
$+\sum_{k^{\prime}=l^{\prime}}^{n-1}\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}$
$=:$ $\sum_{\iota=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\iota}$
@@
$b_{k,l,l^{\prime}}:=$ $\sum_{k^{\prime}=l^{\prime}}^{n-1}\mathbf{1}_{k^{\prime}=k}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j));$
@@
$\tilde{\nu}_{k,l,l^{\prime}}:=$ $\sum_{k=l}^{n-1}\chi_{k}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k^{\prime}+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\mathbf{1}_{k\leq k^{\prime}-1}$
Now note that $b_{k,l,l^{\prime}}\leq\mathrm{Leb}(A)$ and $\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{2}]\lesssim\Delta_{n}$. Furthermore
@@
$\lesssim$ $\Delta_{n}\int_{0}^{n\Delta_{n}}\left(\int_{s}^{\infty}a(u)\mathrm{d}u\right)^{2}\mathrm{d}s$
$\lesssim$ $\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}),$
where in the last inequality we also applied Assumption 1. Similarly, we deduce that that $\mathbb{E}(\rvert\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}\rvert^{2})\lesssim\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1+2(1-\alpha)}).$ In view that $\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]$, $\chi_{k}\nu_{k,l^{\prime},l}$ and $\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}$ are all martingale differences, it follows that
Likewise, for $\iota=2,3$ we have that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}).$
Therefore, in this situation (68) is satisfied with $\upsilon_{n}=C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}.$
@@
$\mathcal{V}_{n}^{(3,3)}$ $=\mathbb{E}(\mathcal{V}_{n}^{(3,3)})+\Delta_{n}^{2}\sum_{l=0}^{n-1}(\overline{c}_{l}^{n})^{2}\mathcal{S}_{l,l}^{n}$
$+2\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{l-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$
@@
The goal of this part is to show that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n}\rvert^{2})\lesssim n\Delta_{n}(l\Delta_{n}+\mathrm{Leb}(A)).$ (70)
The $\sigma$-additivity of $L$ implies that
$\beta_{k,j-1,l}^{(2)}=\sum_{m=k-l}^{j-1}\beta_{k,m,l}^{(1)}.$ (71)
Applying this and performing several tedious changes of order of summation we conclude that $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\imath=1}^{7}\mathcal{S}_{l,l^{\prime}}^{n,\imath}$, where
$\mathcal{S}_{l,l^{\prime}}^{n,1}:=$ $\sum_{m=0}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}[\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}-\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)})]\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
@@
We will only check that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ and $\mathcal{S}_{l,l^{\prime}}^{n,2}$ satisfy (70) since the other terms can be analyzed in the same way. Set $\theta_{k}:=\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$ and note that by the independent scattered property of $L$ it holds that
@@
Lemma 1 implies now that
$\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}\beta_{p,m,l}^{(1)}\beta_{p,m,l^{\prime}}^{(1)})\lesssim_{a,\mathfrak{t}_{L}}\mathrm{Leb}(\cup_{i=0}^{k\land p-l}\mathcal{P}_{A}^{n}(i,m)).$
This estimate along with the fact that $\rvert\theta_{k}\rvert\leq\int_{t_{k}}^{t_{k+1}}a(s)\mathrm{d}s\leq a(0)\Delta_{n}$ lead to
@@
where in the second inequality we also used (21). This shows that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies (70). Similar arguments give that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,2}\rvert^{2})=$ $\sum_{m=l-l^{\prime}}^{n-1}\sum_{m^{\prime}=(l-l^{\prime})\lor(m-l^{\prime})+1}^{m-1}\sum_{k,p=l\lor m+1}^{(n-1)\land(m^{\prime}+l^{\prime})}\mathrm{Leb}(\cup_{i=0}^{k\land p-l}\mathcal{P}_{A}^{n}(i,m))\mathrm{Leb}(\cup_{i=0}^{k\land p-l^{\prime}}\mathcal{P}_{A}^{n}(i,m^{\prime}))\theta_{k}\theta_{p}$
$\lesssim$ $\Delta_{n}\sum_{m=l-l^{\prime}}^{n-1}\sum_{m^{\prime}=(l-l^{\prime})\lor(m-l^{\prime})+1}^{m-1}\sum_{k=l\lor m+1}^{(n-1)\land(m^{\prime}+l^{\prime})}\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,m^{\prime}))\theta_{k}$
@@
Recall that $\varsigma_{k,j}:=\sum_{m=k+1}^{j-1}\alpha_{k+1,m}$ and note that as above we may write
$\mathcal{V}_{n}^{(4,4)}$ $=\mathbb{E}(\mathcal{V}_{n}^{(4,4)})+\Delta_{n}^{2}\sum_{l=0}^{n-1}(\overline{c}_{l}^{n})^{2}\mathcal{S}_{l,l}^{n}$
$+2\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{l-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$
@@
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $\sum_{k=l}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{n-1}\sum_{j=k\lor k^{\prime}+1}^{n}[\varsigma_{k^{\prime},j}\varsigma_{k,j}-\mathbb{E}(\varsigma_{k^{\prime},j}\varsigma_{k,j})]\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$=$ $\sum_{k=l}^{n-1}\sum_{j=k+1}^{n}[\varsigma_{k,j}^{2}-\mathbb{E}(\varsigma_{k,j})]\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$+\sum_{k=l+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-1}\sum_{j=k+1}^{n}\varsigma_{k^{\prime},j}\varsigma_{k,j}\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$+\sum_{k^{\prime}=l+1}^{n-1}\sum_{k=l}^{k^{\prime}-1}\sum_{j=k^{\prime}+1}^{n}\varsigma_{k^{\prime},j}\varsigma_{k,j}\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$=:$ $\sum_{\imath=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$
wherein we have let $\theta_{k,l,k^{\prime},l^{\prime}}^{j}:=\mathrm{Leb}(\cup_{i=0}^{k-l\land k^{\prime}-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$. In this case we are going to show that (68) is satisfied with $\upsilon_{n}=C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}.$ By construction $\varsigma_{k,j}$ is $\mathscr{F}_{k+1}^{n}$-measurable and independent of $\mathscr{F}_{k}^{n}$, reason why $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is a sum of martingale difference for $\imath=1,2,3$. This, along with the fact that $\mathbb{E}(\varsigma_{k,j}^{2})\leq a(0)\Delta_{n}$ and $\sum_{j=k\lor k^{\prime}+1}^{n}\theta_{k,l,k^{\prime},l^{\prime}}^{j}\leq\int_{\rvert t_{k}-t_{k^{\prime}}\rvert}^{\infty}a(s)\mathrm{d}s$ imply that
@@
We now can follow the same arguments as for the case $\ell=2$.
@@
As it is customary, we set
$\mathcal{S}_{l,l^{\prime}}^{n}$ $=\mathfrak{K}_{3}\sum_{k=l^{\prime}}^{n-1}\chi_{k}\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))+\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\chi_{k}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\chi_{k})]\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}]$

### v3

@@ Appendix A Approximations of the Asymptotic Variance
$+\sum_{k=l^{\prime}}^{j-1}\chi_{k}\mathbb{E}(\beta_{k,j,l^{\prime}}^{(1)}\beta_{j,j,l}^{(1)}\chi_{j});$
@@
We now separately analyse the terms
@@
We repeatedly use that $\Delta_{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert$ is uniformly bounded, as shown in the proof of Lemma 6. Our aim is to verify that
@@
where $\mathcal{S}_{l,l^{\prime}}^{n}$ can be written as
$\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\iota=1}^{N}\mathcal{S}_{l,l^{\prime}}^{n,\iota},\qquad N\in\mathbb{N}\text{ fixed},$
and each summand satisfies either
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}\upsilon_{n},\quad\upsilon_{n}=\mathrm{o}(n\Delta_{n}),$ (68)
uniformly on $l,l^{\prime}$, or
@@
for some constant $C\geq 0$. These bounds are sufficient for our purposes. Indeed, by Jensen’s inequality
@@
$\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert\rvert\overline{c}_{l^{\prime}}^{n}\rvert\frac{1}{(n\Delta_{n})^{2}}\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim\frac{1}{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}+1/(n\Delta_{n}).$
In the latter case, as $n\rightarrow\infty$
@@
    If $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1},$ by the continuity of $g^{\prime}$
    $\frac{1}{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}\lesssim T/n\Delta_{n}\rightarrow 0.$
@@
    $\frac{1}{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}\lesssim\mathrm{O}(1/(n\Delta_{n}))+\frac{m_{n}}{n}\times\frac{1}{m_{n}\Delta_{n}}\int_{0}^{m_{n}\Delta_{n}}a(s)^{p+1}s\mathrm{d}s\rightarrow 0,$
@@
Therefore, in any situation, it holds that $\mathbb{E}(\rvert\mathcal{V}_{n}^{(\ell,\ell^{\prime})}-\mathbb{E}(\mathcal{V}_{n}^{(\ell,\ell^{\prime})})\rvert^{2})\rightarrow 0$, as $n\rightarrow\infty$, which is exactly the conclusion of this lemma. In what is left of the proof, for each $\mathcal{V}_{n}^{(\ell,\ell^{\prime})}$ we identify a decomposition of the form (67). Then, we verify that the corresponding summands satisfy either (68) or (69).
For convenience, when $\ell=\ell^{\prime}$ we refer to $\mathcal{V}_{n}^{(\ell,\ell^{\prime})}$ as the conditional variance, and otherwise as the conditional covariance. We also stress that the notation $\mathcal{S}_{l,l^{\prime}}^{n}$ will be reused from case to case: only the decomposition $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\iota=1}^{N}\mathcal{S}_{l,l^{\prime}}^{n,\iota}$ is common, while the precise definition of $\mathcal{S}_{l,l^{\prime}}^{n,\iota}$ depends on the term under consideration.
@@
$\mathcal{S}_{l,l^{\prime}}^{n}:=$ $\mathfrak{K}_{3}\int_{t_{l}}^{t_{l+1}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l^{\prime}}^{(2)}+\mathfrak{K}_{3}\int_{t_{l}}^{t_{l+1}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l}^{(2)}$
$+\int_{0}^{\Delta_{n}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}[\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)})]=:\sum_{\iota=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\iota}.$
Since the sequence $\beta_{j,j-1,l^{\prime}}^{(2)}$ is $l^{\prime}$-dependent and $\mathbb{E}[(\beta_{j,j-1,l^{\prime}}^{(2)})^{2}]\leq\mathrm{Leb}(A),$ it holds that
$\mathbb{E}\left[\left(\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l^{\prime}}^{(2)}\right)^{2}\right]\leq Cn(l^{\prime}+1).$
Hence, $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}(l^{\prime}\Delta_{n}+C).$ The same argument applies to $\mathcal{S}_{l,l^{\prime}}^{n,2}$. Moreover, since $\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}$ is $l\lor l^{\prime}$-dependent, the same bound also holds for $\mathcal{S}_{l,l^{\prime}}^{n,3}$. This is exactly (69).
@@
$\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(2)}\xi_{j,l^{\prime},n}^{(2)}\mid\mathscr{G}_{j-1}^{n})=\sum_{k=l,k^{\prime}=l^{\prime}}^{n-1}\chi_{k}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=(k\lor k^{\prime})+1}^{n-1}\mathcal{P}_{A}^{n}(i,j)).$
@@
$\mathcal{S}_{l,l^{\prime}}^{n}:=\sum_{k=l\lor l^{\prime}}^{n-1}[\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]b_{k,l,l^{\prime}}+\sum_{k=l}^{n-1}\chi_{k}\nu_{k,l^{\prime},l}+\sum_{k^{\prime}=l^{\prime}}^{n-1}\chi_{k^{\prime}}\tilde{\nu}_{k^{\prime},l,l^{\prime}}=:\sum_{\iota=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\iota},$
@@
$b_{k,l,l^{\prime}}:=$ $\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k-l^{\prime})}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j));$
@@
$\tilde{\nu}_{k^{\prime},l,l^{\prime}}:=$ $\sum_{k=l}^{n-1}\chi_{k}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k^{\prime}+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\mathbf{1}_{k\leq k^{\prime}-1}$
Note that $b_{k,l,l^{\prime}}\leq\mathrm{Leb}(A)$ and $\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{2})\lesssim\Delta_{n}$. Hence,
Furthermore
@@
$\lesssim$ $\Delta_{n}\int_{0}^{n\Delta_{n}}\left(\int_{s}^{\infty}a(u)\mathrm{d}u\right)^{2}\mathrm{d}s\lesssim\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}),$
where in the last inequality we also applied Assumption 1. The same bound holds for $\mathbb{E}(\rvert\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}\rvert^{2})$. In view that $\chi_{k}\nu_{k,l^{\prime},l}$ and $\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}$ are martingale differences, it follows that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}),\quad\iota=2,3.$
Therefore (68) is satisfied with $\upsilon_{n}=C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}.$
@@
$\mathcal{V}_{n}^{(3,3)}=\mathbb{E}(\mathcal{V}_{n}^{(3,3)})+\Delta_{n}^{2}\sum_{l=0}^{n-1}(\overline{c}_{l}^{n})^{2}\frac{\mathcal{S}_{l,l}^{n}}{n\Delta_{n}}+2\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{l-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$
@@
Using
$\beta_{k,j-1,l}^{(2)}=\sum_{m=k-l}^{j-1}\beta_{k,m,l}^{(1)},$ (70)
and reordering the sums yields the decomposition $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\imath=1}^{7}\mathcal{S}_{l,l^{\prime}}^{n,\imath}$, in which
$\mathcal{S}_{l,l^{\prime}}^{n,1}:=$ $\sum_{m=0}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l^{\prime})}[\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}-\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)})]\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
@@
It remains to verify that each $\mathcal{S}_{l,l^{\prime}}^{n,\iota}$ satisfies (69). We only focus on $\mathcal{S}_{l,l^{\prime}}^{n,1}$ and $\mathcal{S}_{l,l^{\prime}}^{n,2}$ since the other terms can be analysed in the same way. Set $\theta_{k}:=\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$. Then,
@@
By Lemma 1,
$\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}\beta_{p,m,l}^{(1)}\beta_{p,m,l^{\prime}}^{(1)})\lesssim_{a,\mathfrak{t}_{L}}\mathrm{Leb}(\cup_{i=0}^{k\land p-l}\mathcal{P}_{A}^{n}(i,m)),$
and $\rvert\theta_{k}\rvert\leq a(0)\Delta_{n}$. Hence,
@@
where in the second inequality we also used (21). This shows that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies (69). Similarly, arguments give that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,2}\rvert^{2})=\sum_{m=l-l^{\prime}}^{n-1}\sum_{m^{\prime}=(l-l^{\prime})\lor(m-l^{\prime})+1}^{m-1}$ $\sum_{k,p=l\lor m+1}^{(n-1)\land(m^{\prime}+l^{\prime})}\mathrm{Leb}(\cup_{i=0}^{k\land p-l}\mathcal{P}_{A}^{n}(i,m))$
$\times\mathrm{Leb}(\cup_{i=0}^{k\land p-l^{\prime}}\mathcal{P}_{A}^{n}(i,m^{\prime}))\theta_{k}\theta_{p}$
@@
Recall that $\varsigma_{k,j}:=\sum_{m=k+1}^{j-1}\alpha_{k+1,m}$. Arguing as above we may write
$\mathcal{V}_{n}^{(4,4)}=\mathbb{E}(\mathcal{V}_{n}^{(4,4)})+\Delta_{n}^{2}\sum_{l=0}^{n-1}(\overline{c}_{l}^{n})^{2}\frac{\mathcal{S}_{l,l}^{n}}{n\Delta_{n}}+2\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{l-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$
@@
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $\sum_{k=l}^{n-1}\sum_{j=k+1}^{n}[\varsigma_{k,j}^{2}-\mathbb{E}(\varsigma_{k,j}^{2})]\theta_{k,l,k,l^{\prime}}^{j}+\sum_{k=l+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-1}\sum_{j=k+1}^{n}\varsigma_{k^{\prime},j}\varsigma_{k,j}\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$+\sum_{k^{\prime}=l+1}^{n-1}\sum_{k=l}^{k^{\prime}-1}\sum_{j=k^{\prime}+1}^{n}\varsigma_{k^{\prime},j}\varsigma_{k,j}\theta_{k,l,k^{\prime},l^{\prime}}^{j}=:\sum_{\imath=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$
wherein we have let
$\theta_{k,l,k^{\prime},l^{\prime}}^{j}:=\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l)^{\prime}}\mathcal{P}_{A}^{n}(i,j)).$ (71)
In this case we show that (68) holds with $\upsilon_{n}=C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}.$ By construction $\varsigma_{k,j}$ is $\mathscr{F}_{k+1}^{n}$-measurable and independent of $\mathscr{F}_{k}^{n}$, reason why each $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is a sum of martingale differences. This, along with the fact that $\sum_{j=k\lor k^{\prime}+1}^{n}\theta_{k,l,k^{\prime},l^{\prime}}^{j}\leq\int_{\rvert t_{k}-t_{k^{\prime}}\rvert}^{\infty}a(s)\mathrm{d}s$ imply that
@@
We now can follow the same arguments as in the case $\ell=2$.
@@
Set
$\mathcal{S}_{l,l^{\prime}}^{n}$ $=\mathfrak{K}_{3}\sum_{k=l^{\prime}}^{n-1}\chi_{k}\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))$
$+\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\chi_{k}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\chi_{k})]\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}]$